OpenStudy (anonymous):

Where is the vertex of y = |3x - 6| + 3?

OpenStudy (anonymous):

y=3x-6+3 y=3x-3 find y when x=0 y=3(0)-3 y-3 find x when y=0 0=3x-3 3=3x x=1 vertex is (1,-3)

OpenStudy (anonymous):

but do you write it as (3, -1)?

OpenStudy (anonymous):

i'm just going to go with it, thanks

OpenStudy (anonymous):

no your x comes before y so u write is as (1, -3)